OpenStudy (anonymous):

When y = 12, x = 16. Find x when y = 24.

OpenStudy (anonymous):

like this? y = kx 12 = k(16) k = 3/4 y = (3/4)x 24 = (3/4)x x = 32

OpenStudy (anonymous):

It depends on type of function y is.